### v2

@@ Abstract.
Accurately understanding numbers from financial reports is fundamental to how markets, regulators, algorithms, and normal people read the economy and the world, yet even with XBRL (eXtensible Business Reporting Language) designed to tag every figure with standardized accounting concepts, mapping thousands of facts to over 10,000 U.S. GAAP concepts remains costly, inconsistent, and error-prone. Existing benchmarks define tagging as flat, single-step, extreme classification over small subsets of US-GAAP concepts, overlooking both the taxonomy’s hierarchical semantics and the structured nature of real tagging, where each fact must be represented as a contextualized multi-field output. These simplifications prevent fair evaluation of large language models (LLMs) under realistic reporting conditions. To address these gaps, we introduce 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, the first comprehensive benchmark for structure-aware and full-scope XBRL tagging, designed to evaluate LLMs’ ability to extract and align financial facts through numerical reasoning and taxonomy alignment across text and tables. We define two subtasks: FinNI for numeric identification, which extracts numerical entities and their types from XBRL reports, and FinCL for concept linking, which maps each extracted entity to the corresponding concept in the full US-GAAP taxonomy. Together, these subtasks produce a structured representation of each financial fact. We evaluate diverse LLMs under zero-shot settings and analyze their performance across both subtasks and overall tagging accuracy. Results show that LLMs generalize well in numeric identification but struggle with fine-grained concept linking, revealing current limitations in structure-aware reasoning for accurate financial disclosure. All code and datasets are available on GitHub [^1] https://github.com/The-FinAI/FinTagging/ and Hugging Face [^2] https://huggingface.co/collections/TheFinAI/fintagging-68270132372c6608ac069bef.
XBRL tagging, Benchmark, Large language model, Information extraction, Information retrieval, Reranking
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Test collectionsccs: Information systems Information extractionccs: Information systems Clustering and classification
1. Introduction
Automated tagging is essential in financial reporting, converting structured content such as tables and text into machine-readable data by linking each number to its meaning. Globally, over 2 million companies publish financial reports disclosing earnings, expenses, and liabilities, which provide essential information for investors, regulators, and analysts. However, inconsistent terminology poses challenges for reliable interpretation. To address this, the eXtensible Business Reporting Language (XBRL) was introduced in 1999 as a global standard for machine-readable financial data (Richards et al., 2006). By applying structured tags, as shown in Figure 1, XBRL transforms fragmented disclosures into a consistent format, enabling seamless data exchange, cross-entity comparisons, and large-scale analysis. However, accurately tagging over 2,000 numerical facts per report to more than 10,000 standardized concepts remains labor-intensive and error-prone. In 2023 alone, over 6,500 tagging errors were identified across 33,000 SEC filings [^3] https://xbrl.us/wp-content/uploads/2023/03/DQC-SECMeetingNotes-20240314.pdf.
@@ 1. Introduction
Figure 1. An example of Financial Tagging in Realistic, where the orange numbers mark facts to be tagged, black dashed arrows show the structured outputs after tagging, and blue dashed arrows denote the US-GAAP taxonomy referenced during tagging.
Existing approaches to automated XBRL tagging rely on task-specific models and pretrained language models (PLMs) (Loukas et al., 2022; Sharma et al., 2023), but they require costly annotations, struggle to generalize to new filings, and perform poorly as taxonomies scale. In contrast, recent large language models (LLMs) show strong zero- and few-shot reasoning without task-specific supervision (Wei et al., 2022; Cao et al., 2025; Wang et al., 2025; Peng et al., 2025), yet their application to XBRL tagging remains underexplored due to two major limitations (Table 1). First, task framing is oversimplified: prior benchmarks like FiNER (Loukas et al., 2022) and FNXL (Sharma et al., 2023) treat tagging as extreme multi-class classification over flat label sets, offering little context for disambiguating fine-grained financial facts. These datasets typically cover only 1k+ concepts, leaving most of the 10k+ US-GAAP taxonomy untested. As shown in Table 8, SOTA LLMs such as DeepSeek-V3 (Liu et al., 2024) and GPT-4o (Hurst et al., 2024) achieve 0.0 precision, recall, and F1 under this setting. Second, structured data is ignored: existing datasets exclude tables, despite their central role in financial reporting. Many key facts appear in structured formats, as shown in recent financial QA benchmarks (Chen et al., 2021; Zhu et al., 2021; Mathew et al., 2021), making this omission a key gap between benchmarks and real-world tagging needs.
@@
Table 1. Detailed comparison of financial NLP benchmarks across task types, sources, and structural capabilities. “Num.Reasoning” indicates the numerical reasoning. “Struct.IE” denotes the structured information extraction.
@@
Benchmark Scenario Data Source Task Modality
#Entity
Label
#Taxonomy
Label
Num.
Reasoning?
Struct.
IE?
Concept
Linking?
FinQA (Chen et al., 2021) Decision making SEC 10-K QA text/table 0 0
ConvFinQA (Chen et al., 2022) Decision making SEC 10-K QA text/table 0 0
TAT-QA (Zhu et al., 2021) Financial analysis Chinese financial reports QA text/table 0 0
@@
DocVQA (Mathew et al., 2021) Enterprise automation Financial document images QA text/image 0 0
@@
FiNER-ORD (Shah et al., 2023) Financial tagging Chinese financial disclosures Classification text 49 0
@@
FinRED (Sharma et al., 2022) Knowledge graph construction English financial news Classification text 38 0
FiNER (Loukas et al., 2022) XBRL tagging SEC 10-K Extreme classification text 0 139 / 17,388
@@
FNXL (Sharma et al., 2023) XBRL tagging SEC 10-K Extreme classification text 0 2,800 / 17,388
FinTagging (ours) Financial & XBRL tagging SEC 10-K IE + Alignment text/table 5 17,388 / 17,388
To address these issues, we present 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, the first LLM-ready benchmark for full-scope, structure-aware XBRL tagging, which challenges models to perform end-to-end fact extraction and taxonomy alignment on corporate filings. As shown in Table 1, unlike prior work (Shah et al., 2023; Loukas et al., 2022; Sharma et al., 2023) that frames tagging as flat multi-class classification and focuses solely on narrative text, 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: requires models to jointly extract financial facts and align them with the US-GAAP taxonomy across both unstructured text and structured tables. This is the first LLM-oriented formulation that scales to the full set of 10k+ taxonomy labels while avoiding the intractability of single-step classification, enabling more fine-grained evaluation of model performance. In collaboration with financial reporting experts, we decompose the task into two components: the first, structured numerical information extraction, assesses a model’s ability to identify key financial facts from both text and tables; the second, fine-grained concept linking, evaluates its ability to map each fact to the correct taxonomy concept among all semantically similar candidates. Unlike prior benchmarks limited to 1,000 frequent concepts, our setting covers the entire taxonomy, exposing the limitations of frequency-biased classification and demanding precise semantic reasoning (Bovee et al., 2005). We construct two evaluation sets: FinNI-eval [^4] https://huggingface.co/datasets/TheFinAI/FinNI-eval-v1 for numerical fact extraction and FinCL-eval [^5] https://huggingface.co/datasets/TheFinAI/FinCL-eval-v1 for concept linking, both derived from real XBRL submissions and annotated with gold-standard mappings. We also introduce a unified evaluation framework that jointly measures extraction accuracy and concept alignment, providing a rigorous zero-shot assessment of LLMs’ capabilities for practical, table-aware XBRL tagging.
We evaluate 13 state-of-the-art LLMs under a zero-shot setting on three fronts: (1) end-to-end macro-F1 over the full 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark, (2) subtask-specific performance on FinNI and FinCL, and (3) ablation of our unified extraction-and-alignment evaluation framework. DeepSeek-V3 (Liu et al., 2024) and GPT-4o (Hurst et al., 2024) achieve the highest macro-F1 scores, indicating that our benchmark design enables strong handling of both frequent and rare financial tags, an improvement over traditional PLM baselines evaluated under token-classification settings. In subtask analyses, these LLMs excel at numerical information extraction (FinNI) but continue to struggle with precise concept linking (FinCL), underscoring the remaining challenge of fine-grained semantic alignment. Crucially, our ablation study shows that, without the joint evaluation framework, even top LLMs produce invalid tagging outputs, highlighting the framework’s essential role in realistic assessment of XBRL tagging readiness.
We conclude our main contributions as follows: (1) We introduce 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, the first LLM-ready benchmark for full-scope, table-aware XBRL tagging with a two-stage pipeline, so that modern LLMs can be probed in a true zero-shot setting. For each stage, we release a new, professionally annotated evaluation set: FinNI-eval and FinCL-eval. (2) We conduct an extensive evaluation of state-of-the-art LLMs in zero-shot settings, systematically analyzing their performance across information extraction and semantic alignment subtasks, as well as overall tagging accuracy. (3) Our experimental results reveal a substantial performance gap between LLMs and the task requirements, particularly in fine-grained semantic alignment. This highlights the limitations of current LLMs in complex financial applications and underscores the need for continued advancements in task-specific adaptation.
2. Related Work
XBRL Tagging Benchmarks: Previous benchmarks have framed XBRL tagging as a large-scale classification problem, primarily focusing on entity recognition. FiNER (Loukas et al., 2022) introduced a dataset of 1.1 million sentences from SEC filings, annotated with 139 XBRL entity types, addressing challenges like context-sensitive numeric entities and financial domain-specific expressions. Building on this, Sharma et al. (Sharma et al., 2023) proposed the Financial Numeric Extreme Labelling (FNXL) dataset, expanding the label space to approximately 2,800 XBRL tags and employing extreme classification techniques to handle the increased scale. While these works emphasize the complexity of financial taxonomies, their flat extreme classification approach overlooks the reasoning and alignment needed for realistic XBRL tagging, limiting their suitability for evaluating large language models.
XBRL Tagging Methods: To improve XBRL tagging accuracy, prior work has incorporated structured knowledge and domain-specific modeling. Saini et al. (Saini et al., 2021) proposed GalaXC, a graph neural network leveraging label hierarchies for improved classification. Ma et al. (Ma et al., 2022) enriched transformer models with taxonomy definitions to disambiguate semantically similar tags. Loukas et al. (Loukas et al., 2022) introduced SECBERT and numeral-aware input transformations to enhance robustness. For extreme multi-label tagging, Sharma et al. (Sharma et al., 2023) applied AttentionXML to focus on context-relevant segments. Wang (Wang, 2023) aligned custom tags with standard taxonomies via semantic similarity using TF-IDF, Word2Vec, and FinBERT. More recently, Han et al. (Han et al., 2024) developed XBRL-Agent, an LLM-based system for extracting insights from filings. These efforts highlight the shift toward taxonomy-aware and LLM-enhanced solutions for XBRL tagging.
Financial Evaluation Benchmarks: Broader financial NLP benchmarks address information extraction, reasoning, and document understanding tasks. FiNER-ORD (Shah et al., 2023) and FinRED (Sharma et al., 2022) focus on entity recognition and relation extraction, respectively. BizBench (Koncel-Kedziorski et al., 2023) assesses the quantitative-reasoning ability of LLMs for both business and finance. Pixiu (Xie et al., 2023) evaluates LLMs across classification, QA, and summarization, while FinQA (Chen et al., 2021) and ConvFinQA (Chen et al., 2022) focus on numerical reasoning over text and tables. TAT-QA (Zhu et al., 2021) and DocVQA (Mathew et al., 2021) target table-text joint understanding. These benchmarks highlight the importance of integrating structured and unstructured information, yet they overlook the taxonomy-driven fact alignment and do not support the structured output required for XBRL tagging.
3. 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:
3.1. Task Formulation

### v3

@@ Abstract
Accurate interpretation of numerical data in financial reports is critical for markets and regulators. Although XBRL (eXtensible Business Reporting Language) provides a standard for tagging financial figures, mapping thousands of facts to over 10k US-GAAP concepts remains costly and error-prone. Existing benchmarks oversimplify this task as flat, single-step classification over small subsets of concepts, ignoring the hierarchical semantics of the taxonomy and the structured nature of financial documents. Consequently, these benchmarks fail to evaluate Large Language Models (LLMs) under realistic reporting conditions. To bridge this gap, we introduce 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, the first comprehensive benchmark for structure-aware and full-scope XBRL tagging. We decompose the complex tagging process into two subtasks: (1) FinNI (Financial Numeric Identification), which extracts entities and types from heterogeneous contexts (text and tables); and (2) FinCL (Financial Concept Linking), which maps extracted entities to the full US-GAAP taxonomy. This two-stage formulation enables a fair assessment of LLMs’ capabilities in numerical reasoning and taxonomy alignment. Evaluating diverse LLMs in zero-shot settings reveals that while models generalize well in extraction, they struggle significantly with fine-grained concept linking, highlighting critical limitations in domain-specific structure-aware reasoning. Code is available at GitHub  [^1] https://github.com/The-FinAI/FinTagging and datasets are available at Hugging Face  [^2] https://huggingface.co/collections/TheFinAI/fintagging.
0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:: Benchmarking LLMs for Extracting and Structuring Financial Information
Yan Wang1, Lingfei Qian1, Xueqing Peng1, Yang Ren1, Yi Han2, Keyi Wang3, Dongji Feng4, Fengran Mo5, Shengyuan Lin6, Qinchuan Zhang7, Kaiwen He3, Chenri Luo3, Jianxing Chen3, Junwei Wu3, Chen Xu1, Ziyang Xu1, Jimin Huang1thanks:  Corresponding author, Guojun Xiong8, Xiao-Yang Liu3, Qianqian Xie1, Jian-Yun Nie5 1The Fin AI, 2Georgia Institute of Technology, 3Columbia University, 4California State University 5University of Montreal, 6Carnegie Mellon University, 7Rensselaer Polytechnic Institute, 8Harvard University Correspondence: jimin.huang@thefin.ai
1 Introduction
Although LLMs excel at question answering qian2025fino1, information extraction balasubramanian2025leveraging, and long-document summarization yuan2025strucsum, their ability to perform financial tagging remains underexplored. Unlike open-ended generation, tagging requires models to commit to a precise semantic interpretation for each reported number, grounding it in dispersed evidence across tables and surrounding text, and aligning it to a standardized concept that downstream systems can interpret consistently. Each year, over 2 million companies publish financial reports disclosing earnings, expenses, and liabilities, providing essential information for investors, regulators, and analysts. However, inconsistent terminology complicates reliable interpretation. XBRL, introduced in 1999, defines the global standard for machine-readable financial reporting richards2006introduction and is now adopted in 65 countries [^3] https://www.xbrl.org/the-standard/what/what-is-xbrl/. In practice, XBRL tagging requires grounding thousands of reported numbers per filing in mixed table and text evidence and aligning them to over 17,000 standardized taxonomy concepts(Figure 1), a workflow that remains largely manual and error-prone [^4] https://xbrl.us/wp-content/uploads/2023/03/DQC-SECMeetingNotes-20240314.pdf.
@@ 1 Introduction
Figure 1: An example of Financial Tagging in realistic scenarios, where the blue numbers mark facts to be tagged.
Nevertheless, current XBRL tagging benchmarks often emphasize what is easiest to annotate rather than what is hardest to understand. Many reduce tagging to classification over a truncated taxonomy and remove table structure that carries substantial semantic evidence in real filings (Table 1). As a result, existing datasets typically cover only a few thousand common concepts instead of the full 17,000+ taxonomy, which limits evaluation on rare but important concepts and makes it difficult to assess generalization under long-tailed label distributions. In addition, prior work commonly formulates tagging as closed-set extreme classification loukas2022finer; sharma2023financial. This framing places heavy weight on recalling label identifiers and becomes brittle when the full taxonomy cannot be represented in-context, increasing the likelihood of invalid or hallucinated tags. Consistent with this concern, our ablations show near-zero performance even for strong LLMs such as DeepSeek-V3 liu2024deepseek and GPT-4o hurst2024gpt under full-taxonomy settings. Finally, existing benchmarks focus primarily on narrative text and omit tables, despite tables being the main carrier of numerical disclosures. Consequently, they do not evaluate the joint reasoning over multi-level tabular structure and surrounding text that has been studied in financial QA chen2021finqa; zhu2021tat and is required for realistic end-to-end tagging.
@@
Table 1: Detailed comparison of financial NLP benchmarks across task types, sources, and structural capabilities. “Num.Reasoning” indicates the numerical reasoning. “Struct.IE” denotes the structured information extraction.
@@
Benchmark Scenario Data Source Task Modality #Entity Label #Taxonomy Label Num. Reasoning? Struct. IE? Concept Linking?
FinQA chen2021finqa Decision making SEC 10-K QA text/table 0 0
ConvFinQA chen2022convfinqa Decision making SEC 10-K QA text/table 0 0
TAT-QA zhu2021tat Financial analysis Chinese financial reports QA text/table 0 0
@@
DocVQA mathew2021docvqa Enterprise automation Financial document images QA text/image 0 0
@@
FiNER-ORD shah2023finer Financial tagging Chinese financial disclosures Classification text 49 0
@@
FinRED sharma2022finred Knowledge graph construction English financial news Classification text 38 0
FiNER loukas2022finer XBRL tagging SEC 10-K Extreme classification text 0 139 / 17,688
@@
FNXL sharma2023financial XBRL tagging SEC 10-K Extreme classification text 0 2,800 / 17,688
FINTAGGING Financial & XBRL tagging SEC 10-K IE + Alignment text/table 5 17,688 / 17,688
To address these issues, we introduce 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, the first LLM-oriented benchmark that evaluates whether models can recover the meaning of financial facts by extracting numerical values from mixed table and text inputs and aligning them to the full US-GAAP taxonomy. Unlike prior work loukas2022finer; sharma2023financial that formulates tagging as flat classification over a truncated label set, 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: requires end-to-end reasoning over both narrative disclosures and tables, and scales evaluation to the complete 17k+ concept space. 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: adopts a two-stage task formulation that separates decisions typically conflated in extreme classification. Financial Numerical Identification (FinNI) asks models to identify numerical facts together with their supporting evidence, while Financial Concept Linking (FinCL) maps each fact to the correct US-GAAP concept by semantic comparison with taxonomy definitions, reframing tagging as alignment rather than label recall. We release two expert-verified evaluation sets, FinNI-eval and FinCL-eval, derived from real-world XBRL filings, along with a unified protocol that evaluates both extraction accuracy and concept-alignment performance.
Across 13 state-of-the-art LLMs, our results show a consistent extraction–linking gap. While top models reliably recover numerical facts from mixed table and text inputs, they struggle to assign the correct US-GAAP concepts, indicating that representations sufficient for extraction do not consistently encode the latent semantic conditions that distinguish closely related concepts, such as scope, aggregation level, and qualifying context. As a result, FinNI performance is strong, whereas FinCL remains below 0.19 even after candidate retrieval, with most errors arising from near-neighbor concepts that share similar surface cues but differ in these underlying conditions. Pipeline controls further clarify this failure mode: accuracy improves when gold entities are provided and improves again when the gold concept is included among candidates, yet it does not converge, showing that errors persist even when evidence is localized and the correct concept is available. The restricted cash example illustrates this pattern, where the model selects a narrower concept triggered by a salient local phrase, despite broader contextual evidence supporting an aggregate concept.
We conclude our main contributions as follows: (1) A New Formulation: We introduce 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, the first benchmark that reframes XBRL tagging into a two-stage “Extract-and-Align” paradigm, enabling full-scope evaluation against 17k+ US-GAAP concepts, a scale previously intractable for zero-shot LLMs. (2) High-Fidelity Resources: We release two expert-verified datasets, FinNI-eval and FinCL-eval, which incorporate complex tabular structures to assess joint reasoning across structured and unstructured financial modalities. (3) Diagnostic Insights: By benchmarking 13 SOTA LLMs, we identify a critical “knowledge-alignment gap”, demonstrating that while LLMs excel at numeric extraction, they require our alignment framework to prevent performance collapse during large-scale taxonomy mapping.
2 Related Work
From XBRL Classification to Alignment. Traditional XBRL tagging has been framed as extreme multi-class classification. Benchmarks such as FiNER loukas2022finer and FNXL sharma2023financial evaluate models on truncated label sets (139–2,800 tags), and methods like GalaXC saini2021galaxc, SECBERT loukas2022finer, and AttentionXML sharma2023financial focus on coping with label sparsity through hierarchical modeling. However, these closed-set formulations overlook the semantic reasoning required to navigate the full 17k+ US-GAAP taxonomy and are poorly matched to LLM capabilities. Recent LLM-based tools such as XBRL-Agent han2024xbrlagent explore filing analysis but do not evaluate end-to-end fact-to-taxonomy alignment. 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: addresses this gap by reframing tagging as a two-stage extract-and-align task.
Financial Document Understanding. In parallel, financial NLP has progressed from entity and relation extraction (e.g., FiNER-ORD shah2023finer, FinRED sharma2022finred) to numerical reasoning over documents. Benchmarks such as FinQA chen2021finqa, TAT-QA zhu2021tat, and DocVQA mathew2021docvqa require joint reasoning over text and tables. Yet these tasks are designed for QA and do not constrain outputs to standardized accounting concepts. Our benchmark instead requires both multi-modal fact extraction and precise alignment to official US-GAAP definitions, focusing on the “last mile” of structured financial reporting.
3 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:
3.1 Task Formulation